OpenStudy (anonymous):

Fred earns $6.50 an hour, plus tips, as a waiter. Suppose he works 26 hours in a week. How much money must he receive in tips to earn at least $229 that week?

OpenStudy (anonymous):

6.5 X 26 = 169 229- 169 = your answer

OpenStudy (anonymous):

26(6.50)=169 (No tips) 229-169=60 Tips Total

OpenStudy (anonymous):

thank you very much